OpenStudy (anonymous):

Find the integral of (x-4)/x^2 dx from 1 to 2. (Answer: ln(2)-2.) ( I got 1/x-4/x^2 and should I integrate that from 1 to 2?)

hartnn (hartnn):

yes, do that integrate 1/x - 4/x^2 first :)

OpenStudy (anonymous):

But I have trouble integrating those.

hartnn (hartnn):

integral of 1/x is standard formula \(\int \dfrac{1}{x}dx = \ln x + c\) for 4/x^2 , 1/x^2 = x^-2 and so you can use the x^n formula

OpenStudy (anonymous):

Thanks. I'll try that.

hartnn (hartnn):

could you solve it ?

OpenStudy (anonymous):

Yes, I finally got it. Thank you so much. It worked.

hartnn (hartnn):

most welcome ^_^

OpenStudy (anonymous):

Thanks for the medal.